OpenStudy (anonymous):

what is 32=7x-3 I'm not sure hot to do it i need help

OpenStudy (anonymous):

32 = 7x - 3 35 = 7x x = 5

OpenStudy (anonymous):

add 3 divide by 7

OpenStudy (anonymous):

how did u get the answer

OpenStudy (anonymous):

Whenever you need to solve a single variable equation, you must isolate the variable. To do that in this case, you take your original equation 32 = 7x -3. Add 3 to both sides to get 35 = 7x. Then you need to remove the coefficient of x. To do this, you must divide both sides by 7 and this gives 5 = x.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

ok thank u decemberchild can u help me with another?

OpenStudy (anonymous):

If you'd like.

OpenStudy (anonymous):

ok 153=c2+9 the 2 is and exponet ok

OpenStudy (anonymous):

\[153=c^2+9\]?

OpenStudy (anonymous):

yes thats the equation

OpenStudy (anonymous):

subtract 9, take the square root. don't forget to write \[\pm\] because there are two solutions

OpenStudy (anonymous):

\[153=c^2+9\] \[153-9=c^2\] \[144=c^2\] \[c=\pm\sqrt{144}\] \[c=\pm12\]

OpenStudy (anonymous):

Oh yeah, he is right, idk how I came up with my answer. lol.

OpenStudy (anonymous):

The steps are the same, but my answer is wrong. I need to go to sleep.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Sorry. Do you understand how to do it now though?

OpenStudy (anonymous):

so explain more how the answer is correct what did u do?

OpenStudy (anonymous):

no i dont understand

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

I deleted my incorrect answer. However, you have 153 = c^2 + 9. You need to isolate your variable or in other words, get the variable on one side of the = by itself. So to do this you need to get rid of the 9, so you subtract 9 on both sides to get 144 = c^2. However, you are not finished because you have c^2. In order to get rid of the exponent, you must do the inverse operation of squaring, which is to take the square root. So you take the square root of both sides, and get c = 12, c = -12. The reason there are two answers is due to the fact that anytime you have an equation where the variable has an even exponent, you will get 2 answers.

OpenStudy (anonymous):

thanks youre so much help =} i need help with another if you're ok with it

OpenStudy (anonymous):

One more is fine, then I gotta go.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x+x+x+2=41

OpenStudy (anonymous):

This one is rather easy once you get the concept. You just need to add the x's together to get 3x + 2 = 41. Then you need to isolate the variable again, so you subtract 2 from both sides to get 3x = 39. Then to remove the coefficient of x or the 3, you divide both sides by 3 to get x = 13.

OpenStudy (anonymous):

ok thank u one more please

OpenStudy (anonymous):

Haha, I am sorry, you can attempt it yourself and post your work if you'dlike. I'd be happy to tell you if you are doing it right or not.

OpenStudy (anonymous):

no ill just ask someone else

OpenStudy (anonymous):

ok. have fun.